Two angles that are suplementary differ by 27 degrees. Find the measure of the smaller of the two angles.
Answer by fractalier(6550)

Call one angle x. The other is 180-x. Then we write
x - (180-x) = 27
x - 180 + x = 27
2x = 207
x = 103.5 degrees
180-x = 76.5 degrees, the smaller one.